OpenStudy (anonymous):

Find an exact value. sin 15°

terenzreignz (terenzreignz):

A similar identity exists here \[\Large \sin(\color{red}\alpha-\color{blue}\beta) = \sin(\color{red}\alpha)\cos(\color{blue}\beta) - \cos(\color{red}\alpha) \sin (\color{blue}\beta)\] And, of course, \[\Large 15^o = \color{red}{45^o}- \color{blue}{30^o}\]